OpenStudy (anonymous):

Use the washer method to find the volume of the solid generated by revolving the region bounded by the parabola y=x^4 and the line y=8x about the y-axis.

OpenStudy (anonymous):

did you get stuck somewhere?

OpenStudy (anonymous):

not really, im pretty sure i got the entire thing wrong

OpenStudy (anonymous):

is it the first graph or the one on the bottom?

OpenStudy (anonymous):

x^4, and 8x

OpenStudy (anonymous):

isnt there more to the problem tho? i think there needs to be limits of integration, integrand, and the correct volume

OpenStudy (anonymous):

We have to integrate with respect to Y since we want to use washer method y from 0 to 16

OpenStudy (anonymous):

y=x^4 y=8x x= y^1/4 x= 1/8

OpenStudy (anonymous):

x=1/8 y

OpenStudy (anonymous):

\[\pi \int _0^{16}\left(\sqrt[4]{y}\right)^2-\left(\frac{y}{8}\right)^2dy\]

OpenStudy (zarkon):

the problem is nicer with the shell method \[\int\limits_{0}^{2}2\pi x(8x-x^4)dx\]

OpenStudy (zarkon):

both give the same answer...so it is a nice way to check yourself.

OpenStudy (anonymous):

but my teacher wants us to use the washer method

OpenStudy (anonymous):

the first one is washer

OpenStudy (zarkon):

yes...the first is washers...I just prefer shells for this problem :)

OpenStudy (zarkon):

I get \(\displaystyle\frac{64\pi}{3}\) doing it either way

OpenStudy (anonymous):

alrite thanks you guys